OpenStudy (anonymous):

3(a-2)+8(a-3)=-8 Help.

OpenStudy (anonymous):

distribute so you get 3a-6+8a-24=-8

OpenStudy (anonymous):

a = 2

OpenStudy (anonymous):

so you add 6 to other side then you get 3a + 8a - 24 = -2

OpenStudy (anonymous):

add 24 so you get 3a+8a=22

OpenStudy (anonymous):

add 3a+8a = 11a so 22/11 = 2 a =2

OpenStudy (anonymous):

do you understand???

OpenStudy (anonymous):

you get it?

OpenStudy (anonymous):

I don't get it

OpenStudy (anonymous):

want me to explain again?

OpenStudy (anonymous):

i have an easier way